OpenStudy (anonymous):

Consider the function f(x)=(x^2+x-12)/(x-3).Describe the graph of this function. Include all discontinuities, intercepts, and the basic shape of the graph.

OpenStudy (anonymous):

Notice that \[\frac{x^2+x-12}{x-3}=\frac{(x-3) (x+4)}{x-3}=x+4 \] Can you finish it?

OpenStudy (anonymous):

disc = -4? x int = (3,-4)? y int = 3? basic shape = ellipse?

OpenStudy (anonymous):

or did i mess it all up.